I need help solving this quadratic equation using the Pythagorean Theorem.
Buried treasure. Ahmed has half of a treasure map,
which indicates that the treasure is buried in the
desert 2x + 6 paces from Castle Rock. Vanessa has the
other half of the map. Her half indicates that to find
the treasure, one must get to Castle Rock, walk x paces to the north, and then walk 2x + 4 paces to the east. If they share their information, then they can find x and save a lot of digging. What is x?

Answer by ewatrrr(24785)

 
Hi,
x^2 + (2x+ 4)^2 = (2x+6)^2
x^2 + 4x^2 + 16x + 16 = 4x^2 + 24x + 36
x^2 -8x- 20 = 0 factor
(x+2)(x-10) = 0 (Tossing out the negative solution for unit measure)
x = 10 paces